# ColPali: Efficient Document Retrieval with Vision Language Models

### v3

@@ Abstract
Documents are visually rich structures that convey information through text, as well as tables, figures, page layouts, or fonts. While modern document retrieval systems exhibit strong performance on query-to-text matching, they struggle to exploit visual cues efficiently, hindering their performance on practical document retrieval applications such as Retrieval Augmented Generation. To benchmark current systems on visually rich document retrieval, we introduce the Visual Document Retrieval Benchmark ViDoRe, composed of various page-level retrieving tasks spanning multiple domains, languages, and settings. The inherent shortcomings of modern systems motivate the introduction of a new retrieval model architecture, ColPali, which leverages the document understanding capabilities of recent Vision Language Models to produce high-quality contextualized embeddings solely from images of document pages. Combined with a late interaction matching mechanism, ColPali largely outperforms modern document retrieval pipelines while being drastically faster and end-to-end trainable. We release all project artifacts at https://huggingface.co/vidore.
ColPali: Efficient Document Retrieval with Vision Language Models
Manuel Fayssethanks: Equal Contribution 1,3 Hugues Sibille1,4  Tony Wu1  Bilel Omrani1 Gautier Viaud1Céline Hudelot3Pierre Colombo2,3 1Illuin Technology  2Equall.ai 3CentraleSupélec, Paris-Saclay  4ETH Zürich manuel.faysse@centralesupelec.fr
@@ 1 Introduction
Figure 1: For each term in a user query, ColPali identifies the most relevant document image patches (highlighted zones) and computes a query-to-page matching score. We can then swiftly retrieve the most relevant documents from a large pre-indexed corpus.
@@
Document Retrieval consists in matching a user query to relevant documents in a given corpus. It is central to many industrial applications, either as a standalone ranking system (search engines) or as part of more complex information extraction or Retrieval Augmented Generation (RAG) pipelines.
Over recent years, pretrained language models have enabled large improvements in text embedding models. In practical industrial settings, however, the main performance bottleneck for efficient document retrieval is not in embedding model performance but in the prior data ingestion pipeline. To index a standard PDF document, many steps are required. First, PDF parsers or Optical Character Recognition (OCR) systems are used to extract words from the pages. Document layout detection models can then be run to segment paragraphs, titles, and other page objects such as tables, figures, and headers. A chunking strategy is then defined to group text passages with some semantical coherence, and modern retrieval setups may even integrate a captioning step to describe visually rich elements in a natural language form, more suitable for embedding models. In our experiments (Table 2), we typically find that optimizing the ingestion pipeline yields much greater performance on visually rich document retrieval than optimizing the text embedding model.
Contribution 1: ViDoRe. In this work, we argue that document retrieval systems should not be evaluated solely on the capabilities of text embedding models (Bajaj et al., 2016; Thakur et al., 2021; Muennighoff et al., 2022), but should also consider the context and visual elements of the documents to be retrieved. To this end, we create and openly release ViDoRe, a comprehensive benchmark to evaluate systems on page-level document retrieval with a wide coverage of domains, visual elements, and languages. ViDoRe targets practical document retrieval settings, in which user queries may require both textual and visual understanding to be correctly matched to relevant documents. We highlight the shortcomings of current text-centric systems in these settings. [^1] The benchmark leaderboard is hosted publicly at https://huggingface.co/spaces/vidore/vidore-leaderboard to encourage further developments.
Contribution 2: ColPali. We propose a novel model architecture and training strategy based on Vision Language Models (VLMs) to efficiently index documents purely from their visual features, allowing for subsequent fast query matching with late interaction mechanisms (Khattab and Zaharia, 2020). Our method, ColPali, outperforms all other retrieval systems on ViDoRe while being fast and end-to-end trainable. We release models and code at https://huggingface.co/vidore.
@@ 2 Problem Formulation & Related Work
Efficient document retrieval systems exhibit joint properties of high retrieval performance (R1), low latency during querying (R2), and high throughput during indexation (R3).
@@ 2.1 Textual Retrieval Methods
Neural Retrievers. In bi-encoder models (Reimers and Gurevych, 2019; Karpukhin et al., 2020; Wang et al., 2022), documents are independently mapped offline to a dense vector space. Queries are embedded online and matched to documents through a fast cosine distance computation. A slower, but slightly more performant alternative, cross-encoder systems (Wang et al., 2020; Cohere, 2024) concatenate query and document as a single input sequence and iteratively attribute matching scores to each possible combination. This enables full attention computation between query and document terms but comes at the cost of computational efficiency, as $|\mathcal{D}|$ encoding passes must be done online.
Multi-Vector retrieval via late interaction. In the late interaction paradigm (Khattab and Zaharia, 2020), an embedding is pre-computed and indexed per document token. At runtime, similarity can be computed with individual query token embeddings. The idea is to benefit from the rich interaction between individual query and document terms while taking advantage of the offline computation and fast query matching enabled by bi-encoders.
Retrieval Evaluation. Although benchmarks and leaderboards have been developed to evaluate text embedding models (Thakur et al., 2021; Muennighoff et al., 2022), as previously stated, much of the performance improvements in industrial use cases of embedding models stem from the prior data ingestion pipeline. While documents often rely on visual elements to more efficiently convey information to human readers, text-only systems barely tap into these visual cues.
To our knowledge, no benchmark evaluates document retrieval methods by considering both textual and visual document features like a human would.
@@ 2.2 Integrating Visual features
Contrastive Vision Language Models. Mapping latent representations of textual content to corresponding representations of visual content has been done by aligning disjoint visual and text encoders through contrastive losses (Radford et al., 2021; Zhai et al., 2023). While some OCR capabilities exist in these models, the visual component is often not optimized for text understanding. The Fine-grained Interactive Language-Image Pre-training (Yao et al., 2021) framework extends the late interaction mechanism to cross-modal vision-language models, relying on max similarity operations between text tokens and image patches.
Visually Rich Document Understanding. To go beyond text, some document-focused models jointly encode text tokens alongside visual or document layout features (Appalaraju et al., 2021; Kim et al., 2021; Huang et al., 2022; Tang et al., 2022). Large Language transformer Models (LLMs) with strong reasoning capabilities have recently been combined with Vision Transformers (ViTs) (Dosovitskiy et al., 2020) to create VLMs (Alayrac et al., 2022; Liu et al., 2023b; Bai et al., 2023; Laurençon et al., 2024) where image patch vectors from contrastively trained ViT models (Zhai et al., 2023) are fed as input embeddings to the language model and concatenated with the text-token embeddings.
PaliGemma. The PaliGemma-3B model (Beyer et al., 2024) extends concepts from Pali3 (Chen et al., 2023), and projects SigLIP-So400m/14 (Alabdulmohsin et al., 2023) patch embeddings into Gemma-2B’s text vector space (Gemma Team et al., 2024). Along with its reasonable size w.r.t. other performant VLMs, an interesting property of PaliGemma’s text model is that it is fine-tuned with full-block attention on the prefix (instruction text and image tokens).
@@ 3.1 Benchmark Design
ViDoRe is designed to comprehensively evaluate retrieval systems on their capacity to match queries to relevant documents at the page level. This benchmark encompasses multiple orthogonal subtasks, with focuses on various modalities - text, figures, infographics, tables; thematic domains - medical, business, scientific, administrative; or languages - English (eng), French (fra).
Dataset # Queries Domain
@@
DocVQA (eng) 500 (500) Industrial
InfoVQA (eng) 500 (500) Infographics
TAT-DQA (eng) 1600 (1600) Varied Modalities
arXiVQA (eng) 500 (500) Scientific Figures
TabFQuAD (fra) 210 (210) Tables
@@
Energy (eng) 100 (1000) Scientific
Government (eng) 100 (1000) Administrative
Healthcare (eng) 100 (1000) Medical
AI (eng) 100 (1000) Scientific
Shift Project (fra) 100 (1000) Environment
Table 1: ViDoRe comprehensively evaluates multimodal retrieval methods. The size of the document corpus is indicated in parentheses.
Academic Tasks. We repurpose widely used visual question-answering benchmarks for retrieval tasks: for each page-question-answer triplet, we use the question as the query, and the associated page as the gold document (Table 1). These academic datasets either focus on single specific modalities (Mathew et al., 2020, 2021; Li et al., 2024) or target more varied visually rich documents (Zhu et al., 2022). Moreover, we consider TabFQuAD, a human-labeled dataset on tables extracted from French industrial PDF documents released with this work. Details can be found in subsection A.1.
Practical tasks. We construct topic-specific retrieval benchmarks spanning multiple domains to go beyond repurposed QA datasets and evaluate retrieval in more realistic industrial situations (e.g. RAG). To achieve this, we collect publicly accessible PDF documents and generate queries pertaining to document pages using Claude-3 Sonnet, a high-quality proprietary vision-language model (Anthropic, 2024). In total, we collect 1,000 document pages per topic, which we associate with 100 queries extensively filtered for quality and relevance by human annotators. The corpus topics are intentionally specific to maximize syntactic proximity between documents, creating challenging retrieval tasks and covering an array of orthogonal domains (Table 1). Query-page pair examples are shown in Appendix E. [^2] Answers are generated alongside queries to (1) ground queries and improve their quality and (2) provide resources to foster future work.
Evaluation Metrics. We evaluate performance on our benchmark (Requirement R1) using standard metrics from the retrieval literature (NDCG, Recall@K, MRR). We report NDCG@5 values as the main performance metric in this work and release the complete sets of results along with the models [^3] https://huggingface.co/vidore. To validate compliance with practical industrial constraints, we also consider query latencies (R2) and indexing throughputs (R3).
@@ 3.2 Assessing Current Systems
Results. From a performance perspective, best results are obtained by combining the Unstructured parser with visual information, either from captioning strategies or by running OCR on the visual elements (Table 2). Little difference is seen between BM25 and BGE-M3 embeddings highlighting the visual information bottleneck. Contrastive VLMs lag behind. Beyond retrieval performance (R1), the indexing latencies (R2) reported in Figure 3 illustrate that PDF parsing pipelines can be very lengthy, especially when incorporating OCR or captioning strategies. Querying latencies at runtime (R3) are very good for all evaluated systems ($\leq 22$ ms on NVIDIA L4) due to fast query encoding and cosine similarity matching.
Figure 3: Offline indexing with ColPali is much simpler and faster compared to standard retrieval methods. Indexing speeds reported are computed on Nvidia L4 GPUs and detailed in subsection B.5.
@@ 4.1 Architecture
Vision-Language Models. Encouraged by their strong document understanding capabilities, we propose adapting recent VLMs for retrieval. The key concept is to leverage the alignment between output embeddings of text and image tokens acquired during multi-modal finetuning. To this extent, we introduce ColPali, a Paligemma-3B extension that is capable of generating ColBERT-style multi-vector representations of text and images (Figure 2). PaliGemma-3B is a strong candidate due to its small size, the many released checkpoints fine-tuned for different image resolutions and tasks, and the promising performances on various document understanding benchmarks. We add a projection layer to map the output language modeling embeddings to a vector space of reduced dimension $D=128$ as used in the ColBERT paper (Khattab and Zaharia, 2020) to keep lightweight bag-of-embedding representations.
@@
Contrastive Loss. The Late Interaction operation is fully differentiable, enabling backpropagation. Let a batch $\left\{q_{k},d_{k}\right\}_{k\in[|1,b|]}$ composed of $b$ query-page pairs, where for all $k\in[|1,b|]$, the document page $d_{k}$ is the document corresponding to query $q_{k}$. Following Khattab and Zaharia (2020), we define our in-batch contrastive loss $\mathcal{L}$ as the softmaxed cross-entropy of the positive scores $s_{k}^{+}=\text{LI}\left(q_{k},d_{k}\right)$ w.r.t. to the maximal negative scores $s_{k}^{-}=\max\limits_{l,l\neq k}\hskip 8.53581pt\text{LI}\left(q_{k},d_{l}\right)$.
@@ 4.2 Model training
Dataset. Our training dataset of 127,460 query-page pairs is comprised of train sets of openly available academic datasets ($63\%$) and a synthetic dataset made up of pages from web-crawled PDF documents and augmented with VLM-generated (Claude-3 Sonnet) pseudo-questions ($37\%$). Our training set is fully English by design, enabling us to study zero-shot generalization to non-English languages [^6] Multilingual data is present in the pretraining corpus of the language model (Gemma-2B) and potentially occurs during PaliGemma-3B’s multimodal training.. We explicitly verify no multi-page PDF document is used both ViDoRe and in the train set to prevent evaluation contamination. A validation set is created with $2\%$ of the samples to tune hyperparameters.
@@
Query Augmentation. As in Khattab and Zaharia (2020), we append 5 <unused0> tokens to the query tokens to serve as a soft, differentiable query expansion or re-weighting mechanism.

### v4

@@ Abstract
Documents are visually rich structures that convey information through text, but also figures, page layouts, tables, or even fonts. Since modern retrieval systems mainly rely on the textual information they extract from document pages to index documents -often through lengthy and brittle processes-, they struggle to exploit key visual cues efficiently. This limits their capabilities in many practical document retrieval applications such as Retrieval Augmented Generation (RAG). To benchmark current systems on visually rich document retrieval, we introduce the Visual Document Retrieval Benchmark ViDoRe, composed of various page-level retrieval tasks spanning multiple domains, languages, and practical settings. The inherent complexity and performance shortcomings of modern systems motivate a new concept; doing document retrieval by directly embedding the images of the document pages. We release ColPali, a Vision Language Model trained to produce high-quality multi-vector embeddings from images of document pages. Combined with a late interaction matching mechanism, ColPali largely outperforms modern document retrieval pipelines while being drastically simpler, faster and end-to-end trainable. We release models, data, code and benchmarks under open licenses at https://huggingface.co/vidore.
@@ 1 Introduction
Document Retrieval consists of matching a user query to relevant documents in a given corpus. It is central to many widespread industrial applications, either as a standalone ranking system (search engines) or as part of more complex information extraction or Retrieval Augmented Generation (RAG) pipelines.
Over recent years, pretrained language models have enabled large improvements in text embedding models. In practical industrial settings, however, the primary performance bottleneck for efficient document retrieval stems not from embedding model performance but from the prior data ingestion pipeline. Indexing a standard PDF document involves several steps. First, PDF parsers or Optical Character Recognition (OCR) systems are used to extract words from the pages. Document layout detection models can then be run to segment paragraphs, titles, and other page objects such as tables, figures, and headers. A chunking strategy is then defined to group text passages with some semantical coherence, and modern retrieval setups may even integrate a captioning step to describe visually rich elements in a natural language form, more suitable for embedding models. In our experiments (Table 2), we typically find that optimizing the ingestion pipeline yields much better performance on visually rich document retrieval than optimizing the text embedding model.
Contribution 1: ViDoRe. In this work, we argue that document retrieval systems should not be evaluated solely on the capabilities of text embedding models (Bajaj et al., 2016; Thakur et al., 2021; Muennighoff et al., 2022), but should also consider the context and visual elements of the documents to be retrieved. To this end, we create and openly release ViDoRe, a comprehensive benchmark to evaluate systems on page-level document retrieval with a wide coverage of domains, visual elements, and languages. ViDoRe addresses practical document retrieval scenarios, where queries often necessitate both textual and visual understanding for accurate document matching. We highlight the shortcomings of current text-centric systems in these settings. [^1] The ViDoRe benchmark leaderboard is hosted publicly at https://huggingface.co/spaces/vidore/vidore-leaderboard to encourage further developments.
Contribution 2: ColPali. We propose a novel concept and model architecture based on Vision Language Models (VLMs) to efficiently index documents purely from their visual features, allowing for subsequent fast query matching with late interaction mechanisms (Khattab & Zaharia, 2020). Our method, ColPali, significantly outperforms all other retrieval systems on ViDoRe while being fast and end-to-end trainable. These results demonstrate the potential and the many benefits of this novel Retrieval in Vision Space concept, which could significantly alter the way document retrieval is approached in the industry moving forward. We release all resources at https://hf.co/vidore.
@@ 2 Problem Formulation & Related Work
Under these industrial constraints, we identify three main properties an efficient document retrieval systems should exhibit: (R1) strong retrieval performance, as measured by standard retrieval metrics; (R2) fast online querying, measured through average latencies; (R3) high throughput corpus indexation, ie. the number of pages that can be embedded in a given timeframe.
@@ 2.1 Textual Retrieval Methods
Neural Retrievers. In bi-encoder models (Reimers & Gurevych, 2019; Karpukhin et al., 2020; Wang et al., 2022), documents are independently mapped offline to a dense vector space. Queries are embedded online and matched to documents through a fast cosine distance computation. A slower, but slightly more performant alternative, cross-encoder systems (Wang et al., 2020; Cohere, 2024) concatenate query and document as a single input sequence and iteratively attribute matching scores to each possible combination. This enables full attention computation between query and document terms but comes at the cost of computational efficiency, as $|\mathcal{D}|$ encoding passes must be done online.
Multi-Vector retrieval via late interaction. In the late interaction paradigm introduced by ColBERT (Khattab & Zaharia, 2020), an embedding is pre-computed and indexed per document token. At runtime, similarity can be computed with individual query token embeddings. The idea is to benefit from the rich interaction between individual query and document terms while taking advantage of the offline computation and fast query matching enabled by bi-encoders. See Appendix E for more details.
Retrieval Evaluation. Although benchmarks and leaderboards have been developed to evaluate text embedding models (Thakur et al., 2021; Muennighoff et al., 2022), much of the performance improvements in industrial use cases of embedding models stem from the prior data ingestion pipeline. While documents often rely on visual elements to more efficiently convey information to human readers, text-only systems barely tap into these visual cues. Other work has also independently studied table or chart retrieval systems through repurposed Question Answering datasets (Zhang et al., 2019; Nowak et al., 2024) but only assessing specialized methods for each task.
To our knowledge, no benchmark evaluates document retrieval systems in practical settings; in an end-to-end manner, across several document types and topics, and by evaluating the use of both textual and visual document features.
@@ 2.2 Integrating Visual features
Contrastive Vision Language Models. Mapping latent representations of textual content to corresponding representations of visual content has been done by aligning disjoint visual and text encoders through contrastive losses (Radford et al., 2021; Zhai et al., 2023). While some OCR capabilities exist in these models, the visual component is often not optimized for text understanding.
The Fine-grained Interactive Language-Image Pre-training (Yao et al., 2021) framework extends the late interaction mechanism to cross-modal Vision Language Models, relying on max similarity operations between text tokens and image patches.
Visually Rich Document Understanding. To go beyond text, some document-focused models jointly encode text tokens alongside visual or document layout features (Appalaraju et al., 2021; Kim et al., 2021; Huang et al., 2022; Tang et al., 2022). Large Language transformer Models (LLMs) with strong reasoning capabilities have recently been combined with Vision Transformers (ViTs) (Dosovitskiy et al., 2020) to create VLMs (Alayrac et al., 2022; Liu et al., 2023; Bai et al., 2023; Laurençon et al., 2024b) where image patch vectors from contrastively trained ViT models (Zhai et al., 2023) are fed as input embeddings to the LLM and concatenated with the text-token embeddings.
PaliGemma. The PaliGemma-3B model (Beyer et al., 2024) extends concepts from Pali3 (Chen et al., 2023), and projects SigLIP-So400m/14 (Alabdulmohsin et al., 2023) patch embeddings into Gemma-2B’s text vector space (Gemma Team et al., 2024). Along with its reasonable size w.r.t. other performant VLMs, an interesting property of PaliGemma’s text model is that it is fine-tuned with full-block attention on the prefix (instruction text and image tokens). See Appendix E for more details.
@@ 3.1 Benchmark Design
ViDoRe is designed to comprehensively evaluate retrieval systems on their capacity to match queries to relevant documents at the page level. This benchmark encompasses multiple orthogonal subtasks, with focuses on various modalities - text, figures, infographics, tables; thematic domains - medical, business, scientific, administrative; or languages - English, French. Tasks also span varying levels of complexity, in order to capture signals from both weaker and stronger systems. As many systems require large amounts of time to index pages (captioning-based approaches can take dozens of seconds per page for instance), we limit the number of candidate documents for each retrieval task in order to evaluate even complex systems in a reasonable timeframe without sacrificing quality. For trainable retrieval systems, we provide a reference training set that can be used to facilitate comparisons.
Dataset Language # Queries # Documents Description
@@
DocVQA English 500 500 Scanned documents from UCSF Industry
InfoVQA English 500 500 Infographics scrapped from the web
TAT-DQA English 1600 1600 High-quality financial reports
arXiVQA English 500 500 Scientific Figures from arXiv
TabFQuAD French 210 210 Tables scrapped from the web
@@
Energy English 100 1000 Documents about energy
Government English 100 1000 Administrative documents
Healthcare English 100 1000 Medical documents
AI English 100 1000 Scientific documents related to AI
Shift Project French 100 1000 Environmental reports
Table 1: ViDoRe comprehensively evaluates multimodal retrieval methods.
Academic Tasks. We repurpose widely used visual question-answering benchmarks for retrieval tasks: for each page-question-answer triplet, we use the question as the query, and the associated page as the gold document (Table 1). These academic datasets either focus on single specific modalities (Mathew et al., 2020; 2021; Li et al., 2024) or target more varied visually rich documents (Zhu et al., 2022). Moreover, we consider TabFQuAD, a human-labeled dataset on tables extracted from French industrial PDF documents released with this work. Details can be found in subsection A.1.
Practical tasks. We construct topic-specific retrieval benchmarks spanning multiple domains to go beyond repurposed QA datasets and evaluate retrieval in more realistic industrial situations (e.g. RAG). To achieve this, we collect publicly accessible PDF documents and generate queries pertaining to document pages using Claude-3 Sonnet, a high-quality proprietary vision-language model (Anthropic, 2024). In total, we collect 1,000 document pages per topic, which we associate with 100 queries extensively filtered for quality and relevance by human annotators. The corpus topics are intentionally specific to maximize syntactic proximity between documents, creating more challenging retrieval tasks and covering an array of orthogonal domains (Table 1).  [^2] Answers are generated alongside queries to (1) ground queries and improve their quality and (2) provide resources to foster future work.
Evaluation Metrics. We evaluate performance on our benchmark (Requirement R1) using standard metrics from the retrieval literature (nDCG, Recall@K, MRR). We report nDCG@5 values as the main performance metric in this work and release the complete sets of results along with the models [^3] https://huggingface.co/vidore. To validate compliance with practical industrial requirements (section 2), we also consider query latencies (R2) and indexing throughputs (R3).
@@ 3.2 Assessing Current Systems
Results. From a performance perspective, best results are obtained by combining the Unstructured parser with visual information, either from captioning strategies or by running OCR on the visual elements (Table 2). Little difference is seen between BM25 and BGE-M3 embeddings highlighting the visual information bottleneck. Contrastive VLMs lag behind. Beyond retrieval performance (R1), the indexing latencies (R2) reported in Figure 2 illustrate that PDF parsing pipelines can be very lengthy, especially when incorporating OCR or captioning strategies. Querying latencies at runtime (R3) are very good for all evaluated systems ($\leq 22$ ms on a NVIDIA L4) due to fast query encoding and cosine similarity matching.
Figure 2: Offline document indexing with ColPali is much simpler and faster compared to standard retrieval methods. The PDF Parser results are obtained following the Unstructured settings with BGE-M3 detailed in subsection 3.2. All indexing speeds are averaged per-page latencies. More details in subsection B.4
@@ 4.1 Architecture
Vision-Language Models. Encouraged by their strong document understanding capabilities, we propose adapting recent VLMs for retrieval. The key concept is to leverage the alignment between output embeddings of text and image tokens acquired during multi-modal fine-tuning. To this extent, we introduce ColPali, a Paligemma-3B extension that is capable of generating ColBERT-style multi-vector representations of text and images (Figure 1). PaliGemma-3B is a strong candidate due to its small size, the many released checkpoints fine-tuned for different image resolutions and tasks, and the promising performances on various document understanding benchmarks. We add a projection layer to map each of the language model’s output token embeddings (whether from text or image tokens) to a vector space of reduced dimension $D=128$ as used in the ColBERT paper (Khattab & Zaharia, 2020) to keep lightweight bag-of-embedding representations.
@@
[…]
$\mathcal{L}=-\frac{1}{b}\sum_{k=1}^{b}\log\left[\frac{\exp\left(s_{k}^{+}\right)}{\exp\left(s_{k}^{+}\right)+\exp\left(s_{k}^{-}\right)}\right]=\frac{1}{b}\sum_{k=1}^{b}\log\left(1+\exp\left(s_{k}^{-}-s_{k}^{+}\right)\right)$ (2)
@@ 4.2 Model training
Dataset. Our training dataset of 118,695 query-page pairs is comprised of train sets of openly available academic datasets ($63\%$) and a synthetic dataset made up of pages from web-crawled PDF documents and augmented with VLM-generated (Claude-3 Sonnet) pseudo-questions ($37\%$). Dataset split details are given in subsection A.3. Our training set is fully English by design, enabling us to study zero-shot generalization to non-English languages [^7] Multilingual data is present in the pretraining corpus of the language model (Gemma-2B) and potentially occurs during PaliGemma-3B’s multimodal training.. We explicitly verify no multi-page PDF document is used both ViDoRe and in the train set to prevent evaluation contamination. A validation set is created with $2\%$ of the samples to tune hyperparameters. We openly release the training dataset [^8] https://huggingface.co/datasets/vidore/colpali_train_set for reproducibility and to encourage further research.
@@
Query Augmentation. As in Khattab & Zaharia (2020), we append 5 <unused0> tokens to the query tokens to serve as a soft, differentiable query expansion or re-weighting mechanism.